OpenStudy (anonymous):

the length of a rectangle is 5 centimeters less than it width. The perimeter of the rectangle is 80cm. What are the dimensions of the rectangles what is the answer?

OpenStudy (michele_laino):

If I call with \(w\) and with\(l\) the width and the length respectively, of such rectangle, then we can write the subsequent algebraic system: \[\left\{ \begin{gathered} l = w - 5 \hfill \\ 2\left( {l + w} \right) = 80 \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

please, solve that system, using the substitution method

OpenStudy (anonymous):

i dont understand it @Michele_Laino

OpenStudy (welshfella):

One thing - the length , by definition, is longer than the width.

OpenStudy (michele_laino):

for example, I substitute the first equation into the second one, so I get this: \[\begin{gathered} 2\left( {w - 5 + w} \right) = 80 \hfill \\ \hfill \\ w - 5 + w = 40 \hfill \\ \hfill \\ 2w - 5 = 40 \hfill \\ \end{gathered} \] what is \(w\)?